Professor Davis counts his midterm as 2/3 of the gradeand his final as 1/3 of the grade. Jason scored only 64 on the midterm. What range of scores on the final exam would put Jason's average between 70 and 79 in inclusive?
THANK U MUCH :)

Answer by checkley71(8403)

(2/3)64+(1/3)x=70 or 42 2/3+x/3=70 or x/3=70-42 2/3 or x/3=27 1/3 or x=82
thus Jason must get at least an 82 on his final exam
proof (2/3)64+(1/3)82=70 or 42 1/3+17 1/3=70 or 70=70